

********************************************************************
UARS WEEKLY STATUS REPORT - 31 DEC 1999
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations. SOLSTICE
operates normally, collecting solar data on all available orbital tracking
periods and stellar calibration observations, whenever possible.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM team wishes everyone Happy Holidays!
Daily activity plans for the PEM instrument were successfully generated
for the first two weeks of 2000. PEM health and safety data monitoring, which
was reconfigured for day light operation, are being reviewed to determine
if they can be modified to account for telemetry dropouts caused by the tape
recorder failure. The PEM:: microVAX needs replacement as it is not Y2K
compliant.
The PEM spectrometers continue to be active under PMON control*. While
under PMON control, the PEM spectrometers computed energy deposition from
electrons and protons measured in-situ by the particle spectrometers.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
The CDHF continues to notify all PI Software Representatives when Data Gaps
occur in the LEVEL0 data.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
21 days GPI (4.05) PROD (CLAES L3B) (432R-419C-4F=9L)
41 days WIN (5.11A) PROD (L1-L3) (520R-230C-8F=282L)
UARS CDHF REHOST:
- SUSIM has successfully validated the output products created on the
Rehost System from the SUSIM V7.1 testing.
- Testing HRDI V4.4 on the Rehost System.
- Testing SOLSTICE V4.62 on the Rehost System.
DAAC DATA
Data created in the month of October 1999 for HALOE, NMC, UKMO, and PEM has
been loaded into the DAAC Transfer Tables. DAAC has been notified.
CDHF pending software upgrades: none
***********
FOT
***********
REPORTING PERIOD
This report covers 18 December, 1999 (Orbit 45185, GMT Day 352) through 24
December 1999 (Orbit 45289, GMT Day 358).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 357/2346 45274 Off
The beta angle ranged between the angles of 26.8 degrees to 33.2 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+1.0 degrees on 06 January 1999 (DOY 006).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 17 December 1999 and 20 December 1999 are provided in Appendix D.
The clock error ranged from -12.3 to 10.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 05
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 13
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 06
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/320 - 2287.496800
% Transponder B: 99/334 - 2287.499196
GROUND SYSTEM OPERATION
Three (6) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/357 show the observatory data loss to be 404 hours, 03
minutes, 40.729 seconds (an increase of 61 hours 55 minutes 32.000 seconds
since the last report period). This is a 0.5586 percent data loss which equals
a 99.4414 percent data capture for the mission.
The increase of 61 hours 55 minutes 32.000 seconds occurred over seven (7)
days:
- 04 hours 34 minutes 58.688 seconds on Day 351 due to a S/C TR problem.
- 05 hours 45 minutes 37.248 seconds on Day 352 due to a S/C TR problem.
- 11 hours 30 minutes 11.072 seconds on Day 353 due to a S/C TR problem.
- 10 hours 01 minutes 24.448 seconds on Day 354 due to a S/C TR problem.
- 09 hours 56 minutes 52.960 seconds on Day 355 due to a S/C TR problem.
- 10 hours 10 minutes 50.176 seconds on Day 356 due to a S/C TR problem.
- 09 hours 55 minutes 37.408 seconds on Day 357 due to a S/C TR problem.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS completed a nominal 0.4 degree roll maneuver to return to normal pointing.
This finished up a request from WINDII for their science observations. This
return maneuver was completed Thursday, December 23, 1999 at 18:26:13 Z.
The FOT investigated a repeating low limit violation of the Communication and
Data Handling current. A low current alert was flaging during random
spacecraft contacts. After careful review it was determined that the low
current limit violation was a result of the new tape recorder operations. The
tape recorder draws more current depending on which state it is in, either
playback, record or standby. More specifically, the tape recorder is in the
standby mode during these low limit violations which consequently draws the
least amount of current. Previously the recorder was either in playback mode
or recorder thus never violating the low current limit.
An action has been generated to lower the CDH current limit set to compensate
for the new tape recorder management.
The FOT developed a tape recorder management plan. This plan will be submitted
for approval next week.
YEAR 2000 OPERABILITY SUPPORT
All areas identified to or by the Flight Operations Team for Y2K testing have
been completed. There are no other areas identified at this time that have not
been Y2K certified.
MEETINGS
The FOT did not attend any meetings during this report period.
FUTURE SIGNIFICANT EVENTS
No specific date was agreed upon for UARS Safehold configuration discussions.
Instead, the Supporting Engineers involved will be contacted every two weeks to
determine if either the Flight Operations Team or Supporting Engineers have any
new developments that should be discussed in a group.
UARS has tested and prepared for the Year 2000 event. Special considerations
have been made in preparation for a non nominal 1999 to 2000 transition.
The next UARS Yaw slew is scheduled for January 6, 2000. This maneuver will be
a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 24 December 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
45198 352/2139 99-421 NBTR NOOPS
45213 353/2142 99-422 NBTR NBTR NOOPS
45228 354/2226 99-423 NBTR NBTR NOOPS
45239 355/1506 99-424 OBC Clk Adj fr 0.00 to 20.6 msec/day
45241 355/1826 99-425 NBTR NBTR RTS-58
45243 355/2149 99-426 NBTR NBTR NOOPS
45258 356/2051 99-427 NBTR NBTR NOOPS
45264 357/0712 99-427 NBTR NBTR NOOPS
45261 357/0310 99-428 OBC Clk Adj fr 20.6 to 0.00 msec/day
45270 357/1651 99-429 OBC Roll Maneuver
45272 357/2053 99-430 NBTR NBTR NOOPS
45274 357/2346 99-431 WINDII Power OFF
45285 358/1736 99-430 NBTR NBTR NOOPS
45286 358/1916 99-432 NBTR NBTR NOOPS
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 24 December 1999
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON R-LO PTS Pulse Bus Current Mon 45242
PWICDH Y-LO CDH CURRENT 45242-46,56
PWICDH Y-LO CDH Current 45269
PWICDH Y-LO CDH CURRENT 45276,7,8
HRPTSPBIMON Y-HI PTS Pulse Bus Current Mon 45282-6
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 24 December 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
45239 355/15:11 (21 Dec 99) from 0.00 to 20.6 msec/day
45261 357/03:13 (23 Dec 99) from 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 20 December 1999
17 Dec 1999 - GMT Day 351 Beta = 26.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- --------- ------ -------- ----- ------ ------
1 -44.8/-67.2 -2.73/-3.68 +0.0/+0.0 16.2 0.000 0.0
2 +39.2/-44.8 +7.09/+5.47 +31.2/-15.2 24.5 1.030 21.1
3 +33.6/-16.8 +5.79/+4.51 +30.8/-16.4 24.3 1.023 21.0
20 Dec 1999 - GMT Day 354 Beta = 31.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ---------- ---------- ----- ------ ------
1 -44.8/-67.2 -3.05/-3.99 +0.0/+0.0 16.2 0.000 0.0
2 +22.4/-44.8 +6.11/+5.15 +30.4/-15.2 24.8 1.030 21.5
3 +44.8/-11.2 +5.15/+3.87 +29.6/-16.0 24.8 1.028 21.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ------ ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 24 December 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-089 FOT 45188 352/0516 14170 Generic Late Acquisition
99-091 FOT 45190 352/0938 21717 Fwd link failure at WS
99-092 FOT 45226 354/1823 N/A AP 8 Crashed
99-093 FOT 45256 356/1911 21745 White Sands S/W prob
99-095 FOT 45273 357/2157 N/A AP2 Crash
99-096 FOT 45287 358/2015 21757 CSC Bad TDRSS Vectors
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-091 FOT 45190 352/0938 21717 Fwd link failure at WS
99-092 FOT 45226 354/1823 N/A AP 8 Crashed
99-093 FOT 45256 356/1911 21745 White Sands S/W prob
99-095 FOT 45273 357/2157 N/A AP2 Crash
99-096 FOT 45287 358/2015 21757 CSC Bad TDRSS Vectors
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
99-089 FOT 45095 346/0123 14170 Generic Late Acquisition
SEVNTFW99/52